The State operates a million dollar lottery that pays its winner 20 equal annual payments. The Interest rate to finance this lottery has been 4%. How many $1 tickets needs to be sold for the lottery to break even?
Answer #1

Assuming first payment made when the lottery tickets was purchased ,equal annual installment will amount to 1000000/20=50000

Present value of the winnings at 4% discounting-

50000+50000/1.04+50000/1.04^2+....50000/1.04^19=50000(1-(1/1.04)^20)/(1-(1/1.04))=706697

Hence, Total no of $1 lottery tickets needed to be sold to break even is 706,697/1=706,697